Which of the following is the domain for the equation y − 4.2 = 3.5(x − 2)?
A x ≤ −2.8
B y ≤ −2.8
C −∞ ≤ x ≤ ∞
D −∞ ≤ y ≤ ∞

Found 3 solutions by Theo, greenestamps, math_tutor2020:
Answer by Theo(13342)

equation is y - 4.2 = 3.5 * (x - 2)
add 4.2 to both sides of the equation to get:
y = 3.5 * (x - 2) + 4.2
i think selection C.
that is minus infinity <= x <= plus infinity.
this means any real value of x can be used as input to the function and that the result will be a real value of y for each and every value of x.

Answer by greenestamps(13200)



It's a linear equation with non-zero slope; the domain is all real numbers.

ANSWER: C


Answer by math_tutor2020(3817)


−∞ ≤ x ≤ ∞ should be −∞ < x < ∞

We cannot reach infinity as it's not a number.
It's a concept of going forever in one direction along the number line.

This is why the domain in interval notation would be (−∞, ∞) rather than [−∞, ∞]